OpenStudy (anonymous):

Giving medal to most helpful answer! I am confused on this one? The volume for a rectangular prism is given by the formula V = l · w · h, where l is the length of the prism, w is the width of the prism, and h is the height of the prism. If the volume of a rectangular prism with a length and width of 5 inches is 200 cubic inches, what is the height of the rectangular prism?

OpenStudy (anonymous):

I know that H = V/L*W but, I don't know what the volume is to get the height?

OpenStudy (jdoe0001):

hmm aren't you missing info there? "rectangular prism with a length and width of 5 inches is 200 cubic inches" ^ ?

OpenStudy (jdoe0001):

but anyhow... if you're given L and W and A all you need to do is solve for W

OpenStudy (jdoe0001):

woops.. I mean.solve for H

OpenStudy (anonymous):

-sigh- I still don't understand, 2.5 would be the length and width for both? but, I still don't understand where the 200 cubic inches come in? I don't know what the volume would be?

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

yeah so you need to solve for H cause the L=5 and w=5 and V=200?

OpenStudy (anonymous):

5* not 2.5

OpenStudy (anonymous):

oh but still you need to solve for H? right?

OpenStudy (anonymous):

Oh! I thought you had to like change the 200 cubic inches. But, I gotcha.

OpenStudy (anonymous):

Yeah.

OpenStudy (jdoe0001):

shoiot wrong post =)

OpenStudy (anonymous):

Let me see if I can get the height.

OpenStudy (anonymous):

The height would be 8, correct?

OpenStudy (jdoe0001):

so.. what' was L again?

OpenStudy (anonymous):

5.

OpenStudy (anonymous):

ugh I got 80 xc

OpenStudy (anonymous):

There isn't 80, though. The only 8 digit there is, is 8. .-.

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

sorry Im bad with this :/ i never really understood this type of stuff

OpenStudy (jdoe0001):

hmm so L is 5 and W is 5 and volume is 200 \(\large { V=lwh\quad \begin{cases} l=5\\ w=5\\ V=200 \end{cases}\quad 200=5\cdot 5\cdot h }\)

OpenStudy (anonymous):

No, it's okay. I believe I understand, thankyou for trying though! <3

OpenStudy (anonymous):

Yes, jdoe.

OpenStudy (anonymous):

ooh XD